OpenStudy (anonymous):

(*Medal & Fan* )On Elm Street, there are seven houses with twenty -two windows each. On Maple Street, there are five stores with six windows each. How many windows are there on the two streets?

OpenStudy (anonymous):

@jagr2713 Can you help?

jagr2713 (jagr2713):

Is there an answer choice?

OpenStudy (anonymous):

yes hold on...

OpenStudy (anonymous):

i'm going to try this one on my own because it's kinda easy then can i tag you to check?

jagr2713 (jagr2713):

Ok

OpenStudy (anonymous):

so it's basicly dividing or adding? or even is it both?

jagr2713 (jagr2713):

well its multiplying then adding so yea do you need help anywhere of the problem?

OpenStudy (anonymous):

i just kinda started and had a question but thank's

jagr2713 (jagr2713):

are you stuck anywhere or u got the answer

OpenStudy (anonymous):

no not so far

jagr2713 (jagr2713):

Ok it says 22 windows *each* in 7hours so what do you think

OpenStudy (anonymous):

x

OpenStudy (anonymous):

22*6

OpenStudy (anonymous):

=132

OpenStudy (anonymous):

thats what i did so far

OpenStudy (anonymous):

divide by two?

jagr2713 (jagr2713):

Nope you messed up in the equation.. is 22*6 correct?

OpenStudy (anonymous):

SEVEN!

jagr2713 (jagr2713):

Good so whats the answer now

jagr2713 (jagr2713):

See these are bad mistakes and can give you a whole different answer and can be costly D: Always double check your work :D

OpenStudy (anonymous):

154

OpenStudy (anonymous):

22*7=154

jagr2713 (jagr2713):

Good now do the same for the next one

OpenStudy (anonymous):

next one?

jagr2713 (jagr2713):

Yes it says 6 and 5 so what do we do?

jagr2713 (jagr2713):

Ok you left D: lets finish We do the same 6*5=30 and then it says how many windows are there on the two streets? So we just add them together 154+30= 184

OpenStudy (anonymous):

A.184

OpenStudy (anonymous):

please close this

OpenStudy (anonymous):

um..?